OpenStudy (anonymous):

There are 5,280 feet in a mile. If Sarah ran 13,200 feet, then how many miles did she run?

OpenStudy (anonymous):

1 1/2 miles 5 1/2 miles 3 1/2 miles 2 1/2 miles

OpenStudy (mathstudent55):

You are given the conversion factor as: 5280 ft = 1 mile Write the conversion factor as a fraction. There are two possibilities shown below. \(\dfrac{5280~ft}{1~mile} = \dfrac{1~mile}{5280~ft} = 1\) Now start with the given number: \(13,200~ft\) You want miles, not feet. If you multiply a quantity in ft by a fraction that is miles over ft, the ft will cancel out, and you are left with miles. That means you need to multiply 13,200 ft by \(\dfrac{1~mile}{5280~ft} \). Also, since both conversion fractions equal 1, multiplying by them will not change the amount, just the units. \(13,200 ~ft \times \dfrac{1~mile}{5280~ft} \) Do the multiplication. What do you get?